OpenStudy (anonymous):

How do i simplify: ( y/2 -1 ) ^2

OpenStudy (mathmale):

I'd combine (y/2 -1) into a single fraction first. Can you do that?

OpenStudy (anonymous):

I've forgotten how to do it. I'm supposed to get: (y^2/4 -y/2 -y/2 +1)

OpenStudy (anonymous):

then simplify it further

OpenStudy (mathmale):

How would you combine 1/2 and 1? How would you combine 50 cents and $1? Same principle.

OpenStudy (solomonzelman):

Couple rules to help you out. \(\large\color{blue}{ \rm (a-b)^2=a^2-2ab-b^2 }\) \(\large\color{blue}{ \rm (~\frac{a}{b}~)^{~2}~=~\frac{a^2}{b^2} }\) \(\large\color{blue}{ \rm (-1)^2=1 }\)